By substituting the actual species from the reaction, the expression becomes

rate=k[IO3−][SO32−][H+]

By what factor will the rate of the reaction change if the pH decreases from 4.50 to 2.00?

Express your answer numerically using two significant figures.
Answer #1

find initial [H+]:
use:
pH = -log [H+]
4.5 = -log [H+]
[H+] = 3.162*10^-5 M

find final [H+]:
use:
pH = -log [H+]
2 = -log [H+]
[H+] = 1*10^-2 M

rate=k[IO3−][SO32−][H+]
rate final / rate initial = [H+] final / [H+] initial
= (1*10^-2)/(3.162*10^-5)
= 316
Answer: rate increases by a factor of 320

Limited answer to 2 significant figures